OpenStudy (gdgutierrez):

Couple of series that I cannot figure out. Please help me... Help this poor sleep deprive girl. 1.) Find the MacLaurin series. f(x) = 1/(3-x^2)

OpenStudy (gdgutierrez):

1.)f(x) = \[\frac{ 1 }{ 1-3x^2 }\] \[ans: \sum_{k=0}^{\infty} (2)^n(x)^2n \] \[-\sqrt(2)/2 <x<\sqrt(2)/2\] 2.) \[\lim_{x \rightarrow 0} \frac{e^x-1-x}{x^2}\] 3)\[\int\limits_{}^{} e^(x^2)\] ans:\[\sum_{n=0}^{\infty} (-1)^n \frac{x^{n+1}}{(n+1)!}\]

OpenStudy (anonymous):

I hate to do this with a youtube video, but this one is probably exactly what you want! http://www.youtube.com/watch?v=9EGDZVcl_g8

OpenStudy (anonymous):

if that doesn't make things a lot better for you, i'll try and walk you through it, but it is bedtime!

OpenStudy (gdgutierrez):

Hopefully it is. I gave up that strategy long ago.. Can't sleep when midterm is in 9 hours.

OpenStudy (gdgutierrez):

Thanks for the video though!

OpenStudy (anonymous):

im on spring break! but mine are all next week

OpenStudy (gdgutierrez):

Answered #2 yes!

OpenStudy (dan815):

do u know taylor series

OpenStudy (dan815):

mclauren is taylor when x=0

OpenStudy (anonymous):

:) good job!

OpenStudy (gdgutierrez):

Yeah, ok but I feel like I get lost in the process.

OpenStudy (anonymous):

I stay lost in all the processes of cal 2 :( i haven't even started series yet. i have just been watching the videos to prepare as much as i can.

OpenStudy (gdgutierrez):

This \[\int\limits_{}^{}[1- \frac{x^2}{1!} + \frac{x^4}{2!}- \frac{x^6}{3!}+....] \] is the same as answer for number 3 right?

OpenStudy (gdgutierrez):

Like it's basically saying the same thing but in summation

OpenStudy (dan815):

the integral here means something else

OpenStudy (dan815):

summation only is right

OpenStudy (gdgutierrez):

Because if not I got this using the integral \[\sum_{n=1}^{\infty} (-1)^n \frac{(x^{2n+1})}{(2n+1)(n!)}\]

OpenStudy (dan815):

wha did u do

OpenStudy (gdgutierrez):

I took the integral by using the series of e^x\[e^x = \sum_{n=0}^{\infty} \frac{x^n}{n!}\]

OpenStudy (dan815):

ok yes

OpenStudy (dan815):

wasn't it e^-x^2

OpenStudy (dan815):

it says e^x^2 is that what u want now

OpenStudy (gdgutierrez):

just plugged in after which gave me \[\sum_{n=0}^{\infty} (-1)^n \frac{(x^{2n})}{n!}\]

OpenStudy (gdgutierrez):

took the integral of that. What I want is the series of e^-x^2 which is what I did by manipulating the e^x series.

OpenStudy (dan815):

ok just saying because the question doesn't have a - in there for me

OpenStudy (dan815):

write a couple exapansion out u will see a simple pattern

OpenStudy (dan815):

since x=0 it keeps simplifying

OpenStudy (gdgutierrez):

Like I said I get lost in the process but I'll try not to lose my way haha

OpenStudy (dan815):

ok well to understand mclaruen series and taylor series

OpenStudy (dan815):

its like approximating the next value given the slope accereleration jerk and so on

OpenStudy (dan815):

if u have velocity u know that x=integral v dt = v*t if u have acceleration x=integral integral a dt = 1/2 a t^2.. ignore other terms jerk x= .... = 1/3! j * t^3

OpenStudy (dan815):

thats how you forming this taylor approximation

OpenStudy (anonymous):

\[ \frac {1}{1-u}= \sum_{n=0}^\infty u^n, \quad -1< u <1\\ \frac {1}{1-3x^2}= \sum_{n=0}^\infty (3 x^2)^n, \quad 3 x^2 <1\\ \]

OpenStudy (gdgutierrez):

oooh!! Okay got it. I did not see that for some reason, maybe because it's 3 in the morning here. Thanks!

OpenStudy (anonymous):

YW

OpenStudy (anonymous):

\[ e^x = 1 + x + \frac {x^2}{2!} + \frac {x^3}{3!} + \cdots \\ e^x - 1 - x =\frac {x^2}{2!} + \frac {x^3}{3!} + \cdots \\ \frac{e^x - 1 - x}{x^2} =\frac {1}{2!} + \frac {x}{3!} + \cdots \\ \lim_{x\to 0}\frac{e^x - 1 - x}{x^2}=\lim_{x\to 0}\frac {1}{2!} + \frac {x}{3!} + \cdots=\frac 12 \\ \]

OpenStudy (anonymous):

\[ e^u=\sum_{n=0}^\infty \frac {u^n}{n!}\\ e^{x^2}=\sum_{n=0}^\infty \frac {(x^2)^n}{n!}=\sum_{n=0}^\infty \frac {x^{2n}}{n!}\\ \int e^{x^2}dx=\sum_{n=0}^\infty \int \frac {x^{2n}}{n!}dx=\sum_{n=0}^\infty \frac {x^{2n+1}}{(2n+1)n!} \\ \]

OpenStudy (gdgutierrez):

Thanks! That was reassuring and helpful! Good night!

OpenStudy (anonymous):

Good morning for me.

OpenStudy (anonymous):

just do the division ... 1/3 + 1/9 x^2 + ... ------------------------- 3-x^2 ) 1 -(1 - 1/3 x^2) -------------- 1/3 x^2 -(1/3 x^2 -1/9 x^4) ------------------ 1/9 x^4 ............ \[\sum\frac{x^{2n}}{3^n}\] maybe ....

OpenStudy (anonymous):

to clean it up some ...\[\sum_0^{\infty}~\frac{x^{2n}}{3^{n+1}}\] but thats just a thought